Grade 11Magical Mathematics[Interesting Approach]

What is the value of 00 ? and explain about that deeply.

Lakshmipathi
9 Years agoGrade 11

2 Answers

Vikas TU
9 Years ago
Take example =>
lim x-> 0 (x)^x = L
take log both sides,
that is to say that
logL = xlogx => logx/(1/x) {now it becomes (infinit/infinity)
solve it by that method now.
Apply L-Hospitalinto the limit => (1/x)/(-1/x^2) =  – x = 0  
Now take anit-log both sides,
L = e^0 = 1
Hence the mimit value is 1.
utpal raj
9 Years ago
According to some Calculus textbooks, 0^0 is an ``indeterminate form''. When evaluating a limit of the form 0^0, then you need to know that limits of that form are called ``indeterminate forms''.
we take L hospital rule to solve these type of problems.
 
lim x-> 0 (x)^x = L
take log both sides,
that is to say that
logL = xlogx => logx/(1/x) {now it becomes (infinit/infinity)
solve it by that method now.
Apply L-Hospitalinto the limit => (1/x)/(-1/x^2) =  – x = 0  
Now take anit-log both sides,
L = e^0 = 1
Hence the limit value is 1.